# PyTorchGeoNodes: Enabling Differentiable Shape Programs for 3D Shape Reconstruction

### v1

@@ 4.1 MCTS for Shape Parameter Search
Optimization step. Previous works already combined MCTS with refinement for applications in tasks in 3D scene understanding [32, 33, 2]. Similarly, at the end of MCTS iterations, we perform a gradient-based optimization to further refine values of continuous shape parameters. This is enabled by our PyTorchGeoNodes framework.
@@ 5 Validation
We validate our PyTorchGeoNodes framework and our optimization method for reconstructing shape parameters on four common furniture categories: Cabinet, Sofa, Chair, and Table, for which we created shape programs in Blender. Here we give an overview, and provide more details in the supplementary material:
  • ’Sofa’ has 18 parameters for creating and transforming legs, base, arms, arm legs, and for the presence of an ’L-shaped’ extension and its dimensions;
  • ’Chair’ has 20 parameters that control legs, leg support, backrest, and arms;
  • ’Cabinet’ has 12 parameters for controlling legs, base, back, dividing boards, and drawers;
  • ’Table’ has 10 parameters that control the shape and scale of the top surface, leg types, and middle board.
Our method jointly estimates the 6D pose and shape of the object. We do not need initialization as we do not directly reconstruct global scale in contrast to state-of-the-art approaches such as [2]: In our approach, the object scale is determined by recovering the shape parameters instead, which we designed to model scales of individual parts in meters. This makes our setting more challenging. The orientation is defined as the rotation around the up-axis as done in to [2], which is sufficient for indoor scenes.
Our experiments are divided into two parts: First, we evaluate our method for shape reconstruction on synthetic images of objects. This allows us to evaluate the capabilities of our method in an optimal setup, for example when the object is never occluded. Second, we validate our method on object shape reconstruction for real world scenes from the ScanNet dataset [10]. ScanNet is a very challenging dataset of indoor environment RGB-D scans: Furniture is often only partially visible or even highly occluded, and depth maps are often incomplete due to missing sensor values—this happens for dark and reflective materials, for example. We use SCANnotateDataset [2], which provides ground truth 3D shapes for the objects in ScanNet, in the form of CAD models from ShapeNet [8].
@@ 5.1 Evaluation Criteria
Evaluation of shape parameters. Each object category has individual semantic properties which are an important aspect of the reconstruction. This is neglected by current state-of-the-art methods. With our approach, we integrate these properties as shape parameters into the reconstruction process. We evaluate these properties by comparing the reconstructed and ground truth parameters for the shapes of the target objects. For the experiments in the synthetic domain, we extract the ground truth during the dataset generation process. For the experiments with ScanNet, we manually annotated all relevant objects with semantic properties. We decided to only annotate discrete parameters, as it is extremely challenging and error-prone for humans to manually annotate continuous 3D parameters from images. Instead, we adapt shape similarity metrics.
Evaluation of shape similarity. Similarly to previous work [12, 38], we use the Chamfer distance to quantify the similarity between sampled points of the ground truth and the predicted object. We uniformly sample 10k points on the surface of the predicted and ground truth meshes and use them to compute the Chamfer distance. This metric is important for our experiments on real data when ground truth for continuous shape parameters is difficult to obtain.
5.2 Implementation Details
Data pre-processing. We assume that for the target object in the RGB-D scan, we have an initial estimate for the 3D center of the target object. Then, we estimate 2D segmentation masks of the object for the objective function in Equation (1).
For the synthetic experiment in Section 5.3, we extract the 3D object center and 2D masks during scene generation. In our experiments on ScanNet, we extract these from the annotations from SCANnotateDataset [2]. In general, state-of-the-art 3D object recognition methods could be used instead [37, 28].
Search details. Based on empirical observations from our experiments on the synthetic data, we set the number of MCTS iterations to $300$, with up to $30$ simulations per iteration. We use Adam [17] to perform gradient descent. To deal with large computation overhead, we perform optimization only when the leaf nodes of the tree are reached in the selection phase of MCTS, and skip optimization during simulation. We provide details in the supplementary material.
5.3 Validation on Synthetic Images
For this experiment, we generated $100$ scenes of rendered images of objects generated by the shape programs per category, by randomly sampling shape parameters. Then, we randomly sample $10000$ target points, and render $10$ target views with interpolated rotations around the up-axis. We fix the translation to the origin. In this setting, we therefore evaluate our approach on high-quality images and with precise ground truth annotations.
Table 1 shows quantitative results for cabinets. We observed similar behavior for other categories in the supplementary. We consider three variants: Our full method, MCTS without refinement (’No Refinement’), and MCTS without refinement and the first term of Equation (3). This term encourages exploitation and is typically required for difficult optimization problems.
The full method recovers both types of parameters very accurately, with a somewhat lower number for the ’Has Legs’ parameter. We observe that fine details such as this are very difficult to reconstruct using our current objective function. When disabling refinement, we notice the accuracy of the estimated continuous parameters significantly degrades. The accuracy of the discrete parameters decreases slightly as well, which confirms that our joint discrete and continuous optimization is critical to recovering these parameters. When disabling the exploitation term of MCTS, the accuracy for the discrete parameters additionally drops, which shows that random simulations are not enough and the guided search in MCTS is important for recovering these parameters.
Parameter Full Method No Refinement
No Refinement
No Exploitation Term
@@ 5.3 Validation on Synthetic Images
Continuous Parameters Mean Absolute Difference to Ground Truth (↓)
Width [cm] 5.65 15.72 14.31
Height [cm] 0.99 4.49 5.31
Depth [cm] 2.42 14.16 13.70
Board Thickness [cm] 2.01 2.14 2.14
Leg Width [cm] 1.78 1.77 1.77
Leg Height [cm] 1.61 1.69 1.73
Leg Depth [cm] 1.70 1.71 1.66
Div. Board Thickness [cm] 0.73 0.88 0.88
Rotation [] 1.09 26.19 28.51
Discrete P. Classification Accuracy (↑)
Has Back [%] 98.07 90.38 88.00
Has Legs [%] 82.00 75.00 62.50
Has Drawers [%] 90.00 76.00 79.16
Number of Div. Boards [%] 90.38 90.38 76.00
Table 1: Accuracy of the retrieved shape parameters for ’Cabinet’ on synthetic scenes. For continuous parameters, we report the mean absolute difference to the ground truth parameters. For discrete parameters, we report the classification accuracy. We provide the results for the three variants described in Section 5.3. MCTS does well to reconstruct the discrete parameters, and gradient descent allows to precisely retrieve continuous parameters. Small details such as legs do not contribute much to the objective function, therefore in general we obtain lower accuracy in such situations.
@@ 5.4 Validation on the ScanNet Dataset
Refer to caption Refer to caption
                    our model            GT model our model      GT model
Figure 5: Qualitative comparisons between our reconstructions and the ground truth from ScannotateDataset [2]. First row: Our approach correctly reconstructs a one-armed one-seater, while the CAD model from ScannotateDataset has two armrests. Second row: The width and length of the L extension, and the back height, are too large in the ground truth. Third row: The ground truth is more complex than the actual object. Fourth row: Our approach accurately models the back and legs of the chair, while the ScannotateDataset annotations provide a chair of unrealistic scale. Our method correctly retrieves the height offset of leg support in this example.
Next, we evaluate our approach on challenging real-world scenarios on a subset of validation scenes from ScanNet [10], which provides RGB-D scans of indoor environments. We select a subset of object instances from ScannotateDataset [2] for which we manually annotated 48 cabinets, 67 sofas, 61 chairs, and 100 tables. We consider only object instances that can be approximated by our shape programs and ignore instances that are not yet included in our framework, e.g. tables with built-in cabinets.
Table 2 evaluates quality for object shape reconstruction for sofa and chair categories, we report the numbers for cabinet and table in the supplementary. In addition to classification accuracy, we measure the quality of our reconstructions by comparing them to the ground truth CAD models from SCANnotateDataset. Similarly to synthetic experiments, we maintain solid classification accuracy, except for shape parameters representing small details. At the same time our method is comparable in quality to the ground truth CAD models and we confirm the benefits of our refinement procedure. We emphasize that our method generates good lightweight reconstructions with a small number of vertices and faces: the average for our reconstructed meshes is 64 and 100 respectively, compared to ground truth meshes with 2760 and 10918. Figure 5 shows some qualitative results.
Classification accuracy [%]  (↑) Chamfer
Legs | Left Arm | Right Arm | Arm Legs | Back | L-shaped | Flipped L | Dist. [cm] (↓)
Sofa Random SP 47.69 49.23 58.46 55.38 55.38 44.61 30.0 18.42
No Refinement 43.07 76.92 66.15 64.61 100.0 95.38 100.0 7.2
Full (Ours) 53.84 81.53 72.30 66.15 100.0 98.46 100.0 7.4
Legs Type | Has Leg Support | Back | Has Arms |
Chair Random SP 45.09 45.94 45.09 50.98 16.22
No Refinement 68.62 89.18 98.03 66.66 12.48
Full (Ours) 68.00 91.89 98.00 68.00 11.11
Table 2: Results of shape reconstruction for ’Sofa’ and ’Chair’ categories on the ScanNet scenes. We report classification accuracy for the discrete shape parameters and the Chamfer distance to the CAD annotations SCANnotateDataset [2]. Row-by-row: randomly selected shape parameters as a basic baseline, our method without PyTorchGeoNodes-enabled refinement, and our full approach. For ’Sofa’, the refinement has a significant influence on classification accuracy, except for the legs which are very small and difficult. The ground truth CAD models do not always perfectly represent the scene, which is reflected by similar Chamfer distance values. For ’Chair’, the classification accuracy is not much affected when removing the refinement, which implies MCTS is an important element. By contrast, the Chamfer distance significantly increases, showing the importance of refinement for retrieving the continuous shape parameters accurately.
6 Discussion and Future Work
Refer to caption Refer to caption
Refer to caption Refer to caption
                    retrieved model     GT model retrieved model GT model
Figure 6: Visualizations of current limitations. First row: Even though our shape program considers legs, such details are usually very small, making the reconstruction very challenging especially given the noisy observations from the ScanNet dataset. The annotations from Scannotate have similar issues. Second row: Our Table shape program does not yet consider cabinets built into tables.
We introduced our PyTorchGeoNodes framework that generates efficient PyTorch code for differentiable rendering of interpretable shape programs. In addition, we provided a method for estimating the shape program’s parameters for an object from multiple views.
As Figure 6 shows, our method does come with some limitations. While on synthetic data, our method does well, noisy observations in real scenarios make it difficult to retrieve small details such as the furniture legs. Another limitation is in the expressiveness of hand-designed shape programs. While better expressiveness can be achieved by improving the designs of shape programs, this requires some expertise.
We will make our code publicly available and hope this will encourage other researchers to develop the field. While shape programs are an attractive representation of objects, there are indeed still many things to develop: They could be extended to consider deformable and articulated objects and to integrate textures. Moreover, as mentioned in the introduction, differentiable rendering of shape programs, as allowed by PyTorchGeoNodes, makes it possible to consider development of deep learning methods for creating shape programs from images in a self-supervised way. In this context, as one of many possible future directions, we propose to investigate whether deep networks can be trained to predict shape parameters from images or 3D point clouds while using our PyTorchGeoNodes to enable gradient propagation from various loss functions.

### v2

@@ 4 Search Algorithm
We use our PyTorchGeoNodes framework to compute ${\text{Sh}}({\cal P})$. Thanks to this framework, PyTorch can compute the derivatives of ${\text{Sh}}({\cal P})$ and thus the derivatives of ${\cal L}$ with respect to the continuous parameters in ${\cal P}$. However, some parameters in ${\cal P}$ are discrete, i.e., integers or booleans. Therefore, we cannot optimize ${\cal L}$ using a simple gradient descent. Also, ${\cal L}$ has no special form that we can exploit to optimize it.
To solve this problem, we propose to resort to genetic algorithms to optimize the parameters and combine it with gradient descent to further refine the continuous parameters.
4.1 Shape Parameter Search
A genetic algorithm is a stochastic optimization algorithm that iteratively applies random manipulations of high-dimensional search parameters to identify the most promising complex search subspaces, subject to an objective function. In contrast to gradient-based techniques, it can optimize discrete and continuous values and the objective function can be very general and non-differentiable.
The algorithm starts by randomly generating instances of parameter values ${\cal P}$, individuals, that form an initial set called population of search parameters. Every iteration has three phases: crossover and mutation for exploring search subspaces, and selection for exploiting the most promising subspaces. Crossover randomly selects pairs of individuals, parents, and mixes them by combining the individual parameters to create a new offspring. Therefore, an offspring lies in the search subspace at intersection of parents’ search spaces. Mutation then randomly alters offsprings’ parameters to further explore this subspace. Selection evaluates the quality of the population, and based on survival of the fittest strategy, it takes the best performing individuals to form the population for the next iteration. Further details and hyper-parameters are given in the supplementary material.
Our adaptation of genetic algorithms to our problem is described in details in the supplementary material. The main points are given below. We generate an initial population by randomly generating instances of all program’s parameters ${\cal P}$ representing individuals in the population. We draw random values from the empirical distributions of the parameters. The crossover operation generates offspring $P_{1,2}$ from a pair of parents ${\cal P}_{1}$, ${\cal P}_{2}$. Mutation varies depending on the parameter type. For discrete parameter $d\in{\cal P}_{D}$, we adapt its value $v_{d}$:
$v_{d}\leftarrow\begin{cases}v_{s}\in_{R}{\cal V}_{d}&\quad\text{with probability }P_{m}\\ v_{d}&\quad\text{otherwise},\end{cases}$ (4)
where operator $\in_{R}$ randomly selects value $v$ from the set of valid discrete values ${\cal V}_{p}$. $P_{m}$ is a hyper-parameter that defines the probability of mutating parameter values. In case of a continuous parameter $c\in{\cal P}_{C}$, we adapt its value $v_{c}$ as:
$v_{c}\leftarrow\begin{cases}v_{s}\in_{R}{\cal V}_{c}+{\cal N}(0,\sigma)&\text{with prob. }P_{m}\cdot P_{m}\\ v_{c}+{\cal N}(0,\sigma)&\text{with prob. }P_{m}\\ v_{c}&\text{otherwise},\end{cases}$ (5)
where ${\cal V}_{c}$ is a discretisation of values for parameter $c$ that we further discuss in the supplementary. ${\cal N}(0,\sigma)$ is normal distribution with mean value set to 00 and standard deviation $\sigma$. Finally, we evaluate offsprings, append them to the current population and we select the best individuals based on a score function, in our case negative of objective in Equation 1, to form the population for the next iteration.
Optimization step. Using PyTorchGeoNodes, during the selection phase of every iteration, we perform a gradient-based optimization to further refine values of continuous shape parameters of individuals before selecting the population for next iteration. After a specified number of iterations the search concludes and we return shape parameters of the best performing individual.
4.2 Integrating Gaussians Splats
Shape programs explicitly capture symmetries and similarities between repeated structures, while this aspect is lacking in many representations for 3D reconstruction. On the other hand, earlier shape programs do not capture well details and appearance.
In this section, we show we enable procedural Gaussian splats by integrating them within shape programs and our PyTorchGeoNodes framework. This allows us to capture fine details of objects while keeping the advantages of shape programs. First, we explain how we augment the primitives with ’Gaussian primitives’ while keeping the graph of PyTorchGeoNodes differentiable, and then, we introduce our render-and-compare objective function.
Refer to caption Refer to caption
Original Mesh Augmented with Gaussians
Figure 4: We integrate Gaussians directly into shape programs and PyTorchGeoNodes, to have the advantages of both procedural models and Gaussian splatting.
Extending PyTorchGeoNodes with Gaussian primitives. To enable integration of our procedural graph to support Gaussian splats, we extend the nodes creating mesh primitives to also create trainable Gaussian parameters. We subsample the primitives meshes into many triangles and create Gaussians from these triangles. We initialize the Gaussians means, scales, and quaternions to fit the mesh triangles as we show in Figure 4. Their colors are initialized based on the closest 3D point in the scan. Their opacities are fixed to a high value. As with other primitives, the graph transforms the Gaussians to form a complete set of Gaussians, some of them sharing parameters due to the cloning in the graph.
Optimizing the Gaussians. We noticed that it is more robust to, rather than jointly, first estimate the shape parameters and then optimize the Gaussians. Given a set of optimal shape parameters ${\cal P}$, we minimize the render-and-compare loss as in other Gaussian splatting works, by rendering the Gaussians ${\cal G}({\cal P})$ into scene views [40, 17, 12]:
${\cal L}_{\text{RGB}}+\lambda_{1}{\cal L}_{\text{D}}+\lambda_{2}{\cal L}_{\cal P}+{\cal L}_{\text{reg}}\>,$ (6)
where ${\cal L}_{\text{RGB}}$ and ${\cal L}_{\text{D}}$ are standard losses on color and depth consistency used in Gaussian splatting literature [40, 7, 36].
${\cal L}_{{\cal P}}={\text{CD}}({\text{Sh}}({\cal P}),{\cal G}({\cal P}))$ measures the consistency between the shape parameters and the Gaussians by relying on the chamfer distance. ${\cal L}_{\text{reg}}$ is a regularizer term:
${\cal L}_{\text{reg}}=$ $\lambda_{3}|M({\cal G}({\cal P}))|_{2}+\lambda_{4}|S({\cal G}({\cal P}))|_{2}+$ (7)
$\lambda_{5}\sum_{G\in{\cal G}({\cal P})}\text{Var}(C(G)|,$ (8)
to enforce mean offsets $M({\cal G}({\cal P}))$ and scales offsets $S({\cal G}({\cal P}))$ of Gaussians to be small and the colors of the Gaussians on the same primitive $C(G)$ to remain close to each other. We set $\lambda_{1}=0.1,\lambda_{2}=0.1,\lambda_{3}=100,\lambda_{4}=10.00,\lambda_{5}=1$ in our experiments. In addition, we observed that optimizing object Gaussians and background Gaussians jointly at the same time, with loss for background locations set to $\lambda_{bgd}({\cal L}_{\text{RGB}}+\lambda_{1}{\cal L}_{\text{D}})$ and $\lambda_{bgd}=0.5$ ensures that object Gaussians remain consistent around edges.
@@ 5 Validation
We validate our PyTorchGeoNodes framework and our optimization method on three common furniture categories: Sofa, with $18$ parameters, Chair with $20$, and Table, with $10$, for which we created shape programs in Blender. We designed our shape programs to capture interpretable object structures that are relevant for objects in ScanNet, allowing unseen-before reasoning about high- level object parameters. More details are provided in the supplementary material.
Our method jointly estimates the pose and shape parameters of the object. The orientation is defined as the rotation around the up-axis as done in to [1], which is sufficient for indoor scenes.
We validate our method on object shape reconstruction for real world scenes from the ScanNet dataset [9]. ScanNet is a challenging dataset of indoor environment RGB-D scans: Furniture is only partially visible and highly occluded, and depth maps are incomplete due to missing sensor values—this happens for dark and reflective materials, for example. We manually annotated $176$ objects for Sofa, Chair, and Table. We also provide additional results on automatically generated synthetic scenes in the supplementary material including an additional object category.
@@ 5.1 Evaluation Criteria
Parameter Model from [27] Model from [27] CD CD Genetic Genetic
[…]
@@
Continuous Parameters Mean Absolute Difference to Ground Truth for Table (↓)
Width [cm] 0.74 0.22 0.21 0.21 0.15 0.08
Height [cm] 0.34 0.14 0.15 0.15 0.12 0.06
Depth [cm] 0.4 0.2 0.22 0.22 0.13 0.08
Mid Leg X Scale 0.19 0.13 0.16 0.14 0.1 0.07
Mid Leg Y Scale 0.16 0.14 0.11 0.11 0.08 0.08
Legs Offset Y 0.35 0.34 0.37 0.37 0.31 0.27
Legs Offset X 0.32 0.32 0.36 0.36 0.15 0.13
Legs Scale Y 0.17 0.15 0.18 0.19 0.13 0.1
Mid Board Z Scale 0.21 0.18 0.18 0.18 0.17 0.14
Cabinet Width Scale 0.15 0.1 0.15 0.11 0.1 0.05
Top Thickness [cm] 0.03 0.03 0.02 0.02 0.02 0.02
Legs Size [cm] 0.03 0.03 0.03 0.03 0.01 0.01
Discrete P. Classification Accuracy for Table (↑)
Top Shape 0.78 0.78 0.72 0.72 0.94 0.94
Legs Type 0.39 0.39 0.6 0.58 0.78 0.78
Has Cabinet Leg 0.71 0.71 0.57 0.55 0.78 0.86
Has Mid Board 0.71 0.71 0.55 0.49 0.76 0.76
Reconstruction quality across all categories (↓)
Chamfer distance [m] 0.52 0.21 0.16 0.16 0.14 0.11
Abs. rotation error [rad] 0.45 0.12 0.13 0.11 0.08 0.06
Table 1: Quantitative results of shape program parameters between different baselines for category Table, and general reconstruction metrics across all categories. We observe that GeoCode [27] performs worst due to the domain gap, but the optimization helps improve the performance. Coordinate Descent (CD) does better but is prone to local minima. Genetic algorithm outperforms other baselines by a significant margin.
Evaluation of shape parameters. Each object category has properties which are an important aspect of the reconstruction. This is neglected by current state-of-the-art methods. Our approach integrates these properties as shape parameters into the reconstruction and we evaluate them by comparing the reconstructed and ground truth parameters of the target objects. For the experiments with ScanNet, we manually annotated all relevant objects with shape parameters.
Data pre-processing. We assume that for the target object in the RGB-D scan, we have an initial estimate for the 3D center of the target object, here we follow preprocessing from [1]. Then, we estimate 2D segmentation masks of the object used to generate dense point clouds in the objective function in Equation (1).
5.2 Search Baseline Details
In our evaluations of our search algorithm, we compare results of our genetic algorithm to two baselines:
  • GeoCode [27] introduces a supervised learning method based on an encoder-decoder architecture that takes a target point cloud as input and outputs a value for target program parameters. We use DGCNN [28] as encoder, and the resulting embedding of size $128$ is shared between separate decoders that predict values for program parameters. More precisely, we use three feed-forward layers per parameter, each of size $256$. For discrete parameters, we add a classification head as the output layer with number of outputs matching the number of different discrete values. For continuous parameters, we simply regress the output value. Finally, we perform gradient descent that further refines predictions by passing predicted parameter values to PyTorchGeoNodes and optimizing the objective function. We train it by randomly generating $10000$ instances of ground truth shape program parameters and rotations and then generating synthetic point clouds from these parameters using PyTorchGeoNodes.
  • Our implementation of coordinate descent iteratively optimizes individual parameters of the shape program. First, we initialize the estimate of parameters randomly. Then, at every iteration, we take the current estimate and replace value of the selected parameter by all valid discrete values and calculate the objective term. In case of continuous values we first discretize them as we explain in the supplementary. In case this improves the objective, we update the current estimate. Finally, we perform gradient descent that further refines the current estimate by passing it to PyTorchGeoNodes and optimizing the objective function. We repeat the process for all different parameters and loop 3 times which approximately matches the execution time of our genetic algorithm.
@@ 5.3 Validation on the ScanNet Dataset
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Shape from recovered object parameters aligns very well with the target object also when projected into sampled images of the scene. In the first row, we reconstruct a variety of chairs along with many detailed parameters including seat thickness, thickness and offsets of side and front/back support legs, existence of different semantic elements like ’star-shaped’ legs and its exact rotation. In the second row, we correctly estimate shape of table top and its measurements, including width, depth and thickness, parameters of table legs, and existence of middle shelf including its offset from the bottom. In the third row we reconstruct parameters of ’L-element’ of sofa including its depth and width, existence and size of legs, and existence of individual armrests, backrest, and the corresponding measurements. We show more results in the supplementary material.
We evaluate our approach on challenging real-world scenarios on validation scenes from ScanNet [9], which provides RGB-D scans of indoor scenes. We selected a subset of object instances from the Scannotate dataset [1] for which we manually annotated 57 sofas, 52 chairs, and 67 tables.
Table 1 evaluates the quality for object parameters reconstruction for the ’table’ category and general reconstruction quality across different categories. We report the numbers for ’chair’ and ’sofa’ in the supplementary material. We measure the classification accuracy for discrete parameteres and absolute difference for continuous parameters.
We observe that the supervised learning method from [27], trained on synthetic data, generalises poorly to the scene from ScanNet, while coordinate descent baseline is prone to local minima. Our genetic algorithm performs very well across all metrics except for parameters defining very small details. Visually, the models recovered by our method are comparable in quality to the ground truth and we confirm the benefits of our refinement procedure. Figure 5 and Figure 6 show some qualitative results. In the supplementary we show additional results, and include a downstream application for point cloud segmentation based on our reconstructions.
@@
       (a) (b) (c)
Figure 6: (a) We overlay shape reconstructed using our retrieved shape parameters with the image and (b) compare it to the ground truth. (c) Our integration of Gaussians into PyTorchGeoNodes enables appealing 3D reconstructions.
6 Conclusion
Our code is publicly available and we hope this will encourage other researchers to develop the field and further extend possibilities with PyTorchGeoNodes. While procedural models / shape programs are an attractive representation of objects, there are indeed still many things to develop: e.g. extensions to consider deformable and articulated objects and to integrate textures, or extensions for scene-level procedural models / programs (learned or designed in Blender for example) that would enable reasoning about object relationships and scene physics. Modularity of PyTorchGeoNodes enables integrations with other types of 3D representation, as we demonstrate with our integration of Gaussian Splatting. Moreover, differentiable rendering of PyTorchGeoNodes enables development of deep learning methods for creating shape programs from images in a self-supervised way. Finally, a major limitation is that our shape programs need to be designed by humans for different categories which can lead to limited expressiveness when applying our approach in the wild. Learning shape program designs from database of shapes would make our method easily applicable to novel categories.
Acknowledgement
This project was funded by the European Union (ERC Advanced Grant explorer Funding ID #101097259). This work was performed using HPC resources from GENCI-IDRIS (Grant 2024-AD010615585).